# Worth Their Weight: Randomized and Regularized Block Kaczmarz Algorithms without Preprocessing

### v2

@@ Abstract
Due to the ever growing amounts of data leveraged for machine learning and scientific computing, it is increasingly important to develop algorithms that sample only a small portion of the data at a time. In the case of linear least-squares, the randomized block Kaczmarz method (RBK) is an appealing example of such an algorithm, but its convergence is only understood under sampling distributions that require potentially prohibitively expensive preprocessing steps. To address this limitation, we analyze RBK when the data is sampled uniformly, showing that its iterates converge in a Monte Carlo sense to a weighted least-squares solution. Unfortunately, for general problems the condition number of the weight matrix and the variance of the iterates can become arbitrarily large. We control these issues by incorporating regularization into the RBK iterations, yielding the regularized algorithm ReBlocK. Numerical experiments including examples arising from natural gradient optimization demonstrate that ReBlocK can outperform both RBK and minibatch stochastic gradient descent for inconsistent problems with rapidly decaying singular values.
@@ 1 Introduction
In this work, we are interested in algorithms that solve (1) by sampling just a small number of rows at a time. Such algorithms can be useful for extremely large problems for which direct methods and Krylov subspace methods are prohibitively expensive [6]. Of particular interest are problems in which the dimension $n$ is so large that it is only possible to access $k\ll n$ rows at a time, but not so large that it is necessary to take $k=1$. As a canonical example, we might have $m=10^{9}$, $n=10^{6}$, and $k=10^{3}$.
As further motivation for algorithms of this type, there are some applications in which sampling rows is the only efficient way to access the data. For example, the rows may be computed on the fly, especially when solving the “semi-infinite” version of (1) in which the rows are indexed by continuous variables rather than discrete integers [41]. An example that captures both of these features is the problem of calculating natural gradient directions for continuous function learning problems; see Section 5.
One well-known approach for solving (1) using only a few rows at a time is minibatch stochastic gradient descent (mSGD). The mSGD algorithm works by sampling $k$ rows uniformly at random and using them to calculate an unbiased estimator of the gradient of the least-squares loss function. This is equivalent to averaging the $k$ independent gradient estimators furnished by each sampled row. For a thorough discussion of mSGD, see [23].
The technique of averaging used in mSGD, while simple and efficient, is a relatively crude way of processing a block of $k$ rows. A more sophisticated approach is provided by the randomized block Kaczmarz method, which goes beyond averaging by making use of the pseudoinverse of the sampled block [31]. As we shall see, the use of the pseudoinverse opens up the possibility of faster convergence, but it also creates a number of complications regarding the implementation and analysis of the algorithm.
@@ 1.1 Notation
For symmetric matrices $X,Y$, denote by $X\succ Y$ that the difference $X-Y$ is positive definite, and define ,$\succeq$,$\prec$,$\preceq$ , ≺ , ⪯ correspondingly. For any vector $x$ and any positive definite matrix $Y$ of the same size, let $\left\lVert x\right\rVert_{Y}=\sqrt{x^{\top}Yx}$. For any matrix $X$, denote by $\|X\|_{F}$ its Frobenius norm and by $\sigma_{\rm min}^{+}(X)$ its minimum nonzero singular value.
@@ 1.2 Randomized Kaczmarz
The modern version of the randomized Kaczmarz method (RK) was proposed by Thomas Strohmer and Roman Vershynin in 2009 [42]. In RK, an initial guess $x_{0}$ is updated iteratively using a two-step procedure:
@@
To reduce the sampling cost, it is also possible to run RK with uniform sampling, which is equivalent to running RK on a diagonally reweighted problem [32].
The iteration (3) has the interpretation of projecting $x_{t}$ onto the hyperplane of solutions to the individual equation $a_{i_{t}}^{\top}x=b_{i_{t}}$. For consistent systems, $Ax^{*}=b$, the RK iterates $x_{t}$ converge linearly to $x^{*}$ with a rate that depends on the conditioning of $A$. For inconsistent systems, $Ax^{*}\neq b$, RK converges only to within a finite horizon of the ordinary least-squares solution $x^{*}$ [30]. In particular, the expected squared error $\mathbb{E}\,\left\lVert x_{t}-x^{*}\right\rVert^{2}$ converges to a finite, nonzero value that depends on the conditioning of $A$ and the norm of the residual vector $r=b-Ax^{*}$. See Theorem 7 of [46] for the strongest known convergence bound of this type.
@@ 1.3 Tail Averaging
Tail averaging is a common technique for boosting the accuracy of stochastic algorithms [36, 23, 18]. Given a series of stochastic iterates $x_{0},\ldots,x_{T}$ and a burn-in time $T_{b}$, the tail-averaged estimator is given by
@@
The recent work [18] shows that applying tail averaging to the RK iterates yields exact convergence (with no finite horizon) to the ordinary least-squares solution $x^{*}$, even for inconsistent systems. Building on these results, we will make use of tail averaging to obtain exact convergence to a weighted least-squares solution in the block case.
@@ 1.4 Randomized Block Kaczmarz
Randomized block Kaczmarz (RBK) is an extension of RK which uses blocks of rows to accelerate the convergence and make better use of parallel and distributed computing resources [17, 31]. Like RK, each RBK iteration proceeds in two steps:
@@
Here $A_{S_{t}}^{+}$ is the Moore-Penrose pseudoinverse of $A_{S_{t}}$. The iteration (5) has the interpretation of projecting $x_{t}$ onto the hyperplane of solutions to the block of equations $A_{S_{t}}x=b_{S_{t}}$. The RBK method can also be viewed as a “sketch-and-project” algorithm; see [20].
There have been many proposals for how to choose the blocks in the RBK method. One idea is to use a preprocessing step to partition the matrix $A$ into well-conditioned blocks, then sample this fixed set of blocks uniformly [31]. It has also been suggested to preprocess the matrix with an incoherence transform, which can make it easier to generate a well-conditioned partition [31] or, relatedly, enable RBK with uniform sampling to converge rapidly for the transformed problem [15]. The work of [15] also provides an analysis of the RBK algorithm when sampling from a determinantal point process, and other proposals include greedy block Kaczmarz algorithms such as [26] which require evaluating the complete residual vector at each iteration.
@@
Viewing the problem (1) as a uniform mixture of $m$ distinct rows, we make the natural choice to focus on the case in which the sampling distribution for the algorithm is also uniform. Our results readily generalize to both weighted and semi-infinite problems of the form
@@
for which the corresponding algorithms would independently sample $k$ indices $i_{1},\ldots,i_{k}\sim\mu$. This is especially relevant for scientific applications, in which data is often continuous and may be sampled using a physics-based probability distribution. For example, in the case of neural network wavefunctions [22], the sampling distribution is known as the Born probability density.
@@ 1.5 Contributions
    We provide a new perspective on the pitfalls of RBK with uniform sampling. Concretely, our analysis reveals that when the problem contains many blocks $A_{S}$ that are nearly singular, the condition number of the weight matrix and the variance of the iterates can become arbitrarily large. In Fig. 3 we present numerical examples that manifest these issues.
@@
    We show that RBK-U performs robustly and efficiently for Gaussian data. Concretely, Theorem 3.2 shows that convergence is obtained to the standard least-squares solution when the data arises from a multivariate Gaussian distribution. Furthermore, in this case both the variance of the iterates and the convergence parameter $\alpha$ can be explicitly bounded. When the singular values of the covariance matrix decay rapidly, the convergence rate can be much faster than mSGD; see Corollary 3.3 and Fig. 1.
@@
    We refer to this algorithm as the regularized block Kaczmarz method, or ReBlocK. Similar to RBK-U, we show that ReBlocK with uniform sampling (ReBlocK-U) converges in a Monte Carlo sense to a weighted least-squares solution; see Theorem 4.1. Unlike for RBK-U, both the condition number of the weight matrix and the variance of the iterates can be controlled in terms of just $\lambda$ and some coarse properties of the data $A,b$. This makes ReBlocK-U much more reliable than RBK-U in practice; see Fig. 3. As an added benefit, ReBlocK iterations can be significantly more efficient than RBK iterations; see Sections 4.1 and 8.
@@ 1.6 Related Works
Our regularized algorithm, ReBlocK, is closely related to the iterated Tikhonov-Kaczmarz method [11]. ReBlocK can also be viewed as a specific application of stochastic proximal point algorithms (sPPA) [3, 2, 10] for solving stochastic optimization problems with objective functions of the least-squares type. To ensure the exact convergence of sPPA, diminishing step sizes are required; see for example [34]. In contrast, our work investigates the convergence of sPPA with a large constant step size for the special case of a least-squares loss. Such an approach allows for aggressive updates throughout the algorithm, potentially improving its practical efficiency.
The ReBlocK algorithm is additionally a special case of the SlimLS algorithm proposed by [8] for solving massive linear inverse problems. In fact, their convergence result for the expectation of the ReBlocK iterates, found in result (a) of their Theorem 3.1, is equivalent to equation (24) of our Theorem 4.1. However, by identifying the limiting expectation value as a weighted least-squares solution, we provide greater conceptual clarity regarding its meaning. Additionally, our bounds regarding the variance of the ReBlocK iterates, the convergence of tail averages, and the suboptimality of the weighted least-squares solution are all much stronger than the corresponding bounds of [8]. These stronger bounds serve to explain our numerical findings that ReBlocK can be robust and efficient even when $\lambda$ is very small and the problem contains many nearly singular blocks $A_{S}$, which is the most relevant case and is not explained by the previous analysis.
Our work is also related to the nearly concurrent paper [16], which introduces Tikhonov regularization into the RBK iterations just like ReBlocK. [16] focuses on consistent systems and in this context, the regularization gives rise to optimal convergence rates in the presence of Nesterov acceleration. On the other hand, our work focuses on solving inconsistent systems, in which case the regularization is needed to ensure the stability of the algorithm. Exploring the combination of regularization and Nesterov acceleration in the inconsistent case is a promising direction for future work.
@@ 2 Overview of the Analysis
Here $M(\cdot):\mathbb{R}^{k\times n}\rightarrow\mathbb{R}^{k\times k}$ takes in the sampled block $A_{S_{t}}$ and returns a positive semidefinite “mass” matrix. RBK is recovered by setting $M(A_{S_{t}})=(A_{S_{t}}A_{S_{t}}^{\top})^{+}$ and ReBlocK by setting $M(A_{S_{t}})=(A_{S_{t}}A_{S_{t}}^{\top}+\lambda|S_{t}|I)^{-1}$. See Algorithm 1 for the full procedure with and without tail averaging.
@@
where $I_{S}$ represents the rows of the $n\times n$ identity matrix whose indices are in $S$. These quantities are natural because they enable the general iteration (8) to be rewritten as
@@
Since the normal equations for (12) can be written as $A^{\top}\overline{W}r^{(\rho)}=0$, we observe that the final term in (13) vanishes in expectation. Indeed, identifying the appropriate weighted solution $x^{(\rho)}$ to enable the generalized iteration (8) to be written in the form of (13), namely as a linear contraction of the error plus a zero-mean additive term, is the main technical innovation underlying our results. From here, the analysis of [18] can be readily generalized to show that convergence to $x^{(\rho)}$ is obtained.
@@ 3 RBK without Preprocessing
(15) $\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq\left(1-\alpha\right)^{T_{b}+1}\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}+\frac{1}{\alpha^{2}(T-T_{b})}\;\mathbb{E}\,_{S\sim\rho}\left\lVert A_{S}^{+}r^{(\rho)}_{S}\right\rVert^{2}.$
@@
To our knowledge, this is the first result for inconsistent linear systems that characterizes the exact solution to which the randomized block Kaczmarz iterates converge, albeit in a Monte Carlo sense. The $O(1/T)$ convergence rate for the tail-averaged bound is optimal for row-access methods, and a reasonable default for the burn-in time is $T_{b}=T/2$; see [18] for a more thorough discussion of these points. The proof of Theorem 3.1, which takes advantage of an orthogonal decomposition of the error term $x_{t+1}-x^{(\rho)}$, can be found in Appendix B.
Unfortunately, Theorem 3.1 does not imply robust convergence for general problems. For one, the convergence parameter $\alpha$, which affects every term in the bound, is difficult to analyze in general. Worse, for problems containing nearly singular blocks $A_{S}$, the weight matrix $\overline{W}=\mathbb{E}_{S\sim\rho}\left[I_{S}^{\top}(A_{S}A_{S}^{\top})^{+}I_{S}\right]$ can be arbitrarily poorly conditioned and the variance term $\mathbb{E}\,_{S\sim\rho}||A_{S}^{+}r^{(\rho)}_{S}||^{2}$ can be arbitrarily large. See Fig. 3 for numerical examples that manifest these problems.
In the next subsection, we show that these problems are avoided when the data arises from a Gaussian distribution. A more general solution is provided by the ReBlocK algorithm introduced in Section 4.
3.1 Linear Least-squares with Gaussian Data
Consider a random problem with the data $A,b$ generated as
(16) $\begin{bmatrix}a_{i}^{\top}&b_{i}\end{bmatrix}\sim\mathcal{N}(0,Q)$

### v3

@@ Abstract
Due to the ever growing amounts of data leveraged for machine learning and scientific computing, it is increasingly important to develop algorithms that sample only a small portion of the data at a time. In the case of linear least-squares, the randomized block Kaczmarz method (RBK) is an appealing example of such an algorithm, but its convergence is only understood under sampling distributions that require potentially prohibitively expensive preprocessing steps. To address this limitation, we analyze RBK when the data is sampled uniformly, showing that its iterates converge in a Monte Carlo sense to a weighted least-squares solution. Unfortunately, for general problems the bias of the weighted least-squares solution and the variance of the iterates can become arbitrarily large. We show that these quantities can be rigorously controlled by incorporating regularization into the RBK iterations, yielding the regularized algorithm ReBlocK. Numerical experiments including examples arising from natural gradient optimization demonstrate that ReBlocK can outperform both RBK and minibatch stochastic gradient descent for inconsistent problems with rapidly decaying singular values.
@@ 1 Introduction
In this work, we are interested in algorithms that solve (1) by sampling just a small number of rows at a time. Such algorithms can be useful for extremely large problems for which direct methods and Krylov subspace methods are prohibitively expensive [censor1981row]. Of particular interest are problems in which the dimension $n$ is so large that it is only possible to access $k\ll n$ rows at a time, but not so large that it is necessary to take $k=1$. As a canonical example, we might have $m=10^{9}$, $n=10^{6}$, and $k=10^{3}$.
As further motivation for algorithms of this type, there are some applications in which sampling rows is the only efficient way to access the data. For example, the rows may be computed on the fly, especially when solving the “semi-infinite” version of (1) in which the rows are indexed by continuous variables rather than discrete integers [shustin2022semi]. An example that captures both of these features is the problem of calculating natural gradient directions for continuous function learning problems; see Section 6.
One well-known approach for solving (1) using only a few rows at a time is minibatch stochastic gradient descent (mSGD). The mSGD algorithm works by sampling $k$ rows uniformly at random and using them to calculate an unbiased estimator of the gradient of the least-squares loss function. This is equivalent to averaging the $k$ independent gradient estimators furnished by each sampled row. For a thorough discussion of mSGD, see [jain2018parallelizing].
The technique of averaging used in mSGD, while simple and efficient, is a relatively crude way of processing a block of $k$ rows. A more sophisticated approach is provided by the randomized block Kaczmarz method, which goes beyond averaging by making use of the pseudoinverse of the sampled block [needell2014paved]. As we shall see, the use of the pseudoinverse opens up the possibility of faster convergence, but it also creates a number of complications regarding the implementation and analysis of the algorithm.
@@ 1.1 Notation
For symmetric matrices $X,Y$, denote by $X\succ Y$ that the difference XYX-Y is positive definite, and define $\succeq,\prec,\preceq$ correspondingly. For any vector $x$ and any positive definite matrix $Y$ of the same size, let $\left\lVert x\right\rVert_{Y}=\sqrt{x^{\top}Yx}$. For any matrix $X$, denote by $\|X\|_{F}$ its Frobenius norm and by $\sigma_{\rm min}^{+}(X)$ its minimum nonzero singular value.
@@ 1.2 Randomized Kaczmarz
The modern version of the randomized Kaczmarz method (RK) was proposed by Thomas Strohmer and Roman Vershynin in 2009 [strohmer2009randomized]. In RK, an initial guess $x_{0}$ is updated iteratively using a two-step procedure:
@@
To reduce the sampling cost, it is also possible to run RK with uniform sampling, which is equivalent to running RK on a diagonally reweighted problem [needell2014stochastic].
The iteration (3) has the interpretation of projecting $x_{t}$ onto the hyperplane of solutions to the individual equation $a_{i_{t}}^{\top}x=b_{i_{t}}$. For consistent systems, $Ax^{*}=b$, the RK iterates $x_{t}$ converge linearly to $x^{*}$ with a rate that depends on the conditioning of $A$. For inconsistent systems, $Ax^{*}\neq b$, RK converges only to within a finite horizon of the ordinary least-squares solution $x^{*}$ [needell2010randomized]. In particular, the expected squared error $\mathbb{E}\,\left\lVert x_{t}-x^{*}\right\rVert^{2}$ converges to a finite, nonzero value that depends on the conditioning of $A$ and the norm of the residual vector $r=b-Ax^{*}$. See Theorem 7 of [zouzias2013randomized] for the strongest known convergence bound of this type.
@@ 1.3 Tail Averaging
Tail averaging is a common technique for boosting the accuracy of stochastic algorithms [rakhlin2011making, jain2018parallelizing, epperly2024randomizedkaczmarztailaveraging]. Given a series of stochastic iterates $x_{0},\ldots,x_{T}$ and a burn-in time $T_{b}$, the tail-averaged estimator is given by
@@
The recent work [epperly2024randomizedkaczmarztailaveraging] shows that applying tail averaging to the RK iterates yields exact convergence (with no finite horizon) to the ordinary least-squares solution $x^{*}$, even for inconsistent systems. Building on these results, we will make use of tail averaging to obtain exact convergence to a weighted least-squares solution in the block case.
@@ 1.4 Randomized Block Kaczmarz
Randomized block Kaczmarz (RBK) is an extension of RK which uses blocks of rows to accelerate the convergence and make better use of parallel and distributed computing resources [elfving1980block, needell2014paved]. Like RK, each RBK iteration proceeds in two steps:
@@
Here $A_{S_{t}}^{+}$ is the Moore-Penrose pseudoinverse of $A_{S_{t}}$. The iteration (5) has the interpretation of projecting $x_{t}$ onto the hyperplane of solutions to the block of equations $A_{S_{t}}x=b_{S_{t}}$. The RBK method can also be viewed as a “sketch-and-project” algorithm; see [gower2015randomized].
There have been many proposals for how to choose the blocks in the RBK method. One idea is to use a preprocessing step to partition the matrix $A$ into well-conditioned blocks, then sample this fixed set of blocks uniformly [needell2014paved]. It has also been suggested to preprocess the matrix with an incoherence transform, which can make it easier to generate a well-conditioned partition [needell2014paved] or, relatedly, enable RBK with uniform sampling to converge rapidly for the transformed problem [derezinski2024solving]. The work of [derezinski2024solving] also provides an analysis of the RBK algorithm when sampling from a determinantal point process, and other proposals include greedy block Kaczmarz algorithms such as [liu2021greedy] which require evaluating the complete residual vector at each iteration.
@@
To avoid preprocessing, we consider the case that the sampling strategy is chosen a priori in that it is not necessarily adapted to the data matrix $A$ or the right-hand side $b$. To simplify the analysis, we focus concretely on the case when the sampling is uniform. This is a natural choice since (1) can be viewed as a uniform mixture of $m$ distinct rows. We emphasize that this choice is made for simplicity and that our results are not inherently restricted to the case of uniform sampling. For example, our results can be readily generalized to both weighted and semi-infinite problems of the form
@@
for which the corresponding algorithms would independently sample $k$ indices $i_{1},\ldots,i_{k}\sim\mu$. This is especially relevant for scientific applications, in which data is often continuous and may be sampled using a physics-based probability distribution. For example, in the case of neural network wavefunctions [hermann2023ab], the sampling distribution is known as the Born probability density.
@@ 1.5 Contributions
    We provide a new perspective on the pitfalls of RBK with uniform sampling. Concretely, our analysis reveals that when the problem contains many blocks $A_{S}$ that are nearly singular, the bias of the weighted least-squares solution and the variance of the iterates can become arbitrarily large. See Theorems 3.1 and 18 for theoretical barriers and Fig. 2 for numerical examples that manifest these issues.
@@
    We show that RBK-U performs robustly and efficiently for Gaussian data. Concretely, Theorem 4.1 shows that convergence is obtained to the ordinary least-squares solution when the data arises from a multivariate Gaussian distribution. Furthermore, in this case both the variance of the iterates and the convergence parameter $\alpha$ can be explicitly bounded. When the singular values of the covariance matrix decay rapidly, the convergence rate can be much faster than mSGD; see Corollary 4.2 and Fig. 3.
@@
    We refer to this algorithm as the regularized block Kaczmarz method, or ReBlocK. Similar to RBK-U, we show that ReBlocK with uniform sampling (ReBlocK-U) converges in a Monte Carlo sense to a weighted least-squares solution; see Theorem 5.1. Unlike for RBK-U, the bias of the weighted least-squares solution and the variance of the iterates can be controlled in terms of just $\lambda$ and some coarse properties of the data $A,b$. This makes ReBlocK-U much more reliable than RBK-U in practice; see Figs. 6 and 8. As an added benefit, ReBlocK iterations can be significantly more efficient than RBK iterations; see Sections 5.2 and 12.
@@ 1.6 Related Works
Our regularized algorithm, ReBlocK, is closely related to the iterated Tikhonov-Kaczmarz method [de2011modified]. ReBlocK can also be viewed as a specific application of stochastic proximal point algorithms (sPPA) [bertsekas2011incremental, asi2019stochastic, davis2019stochastic] for solving stochastic optimization problems with objective functions of the least-squares type. To ensure the exact convergence of sPPA, diminishing step sizes are required; see for example [puatracscu2021new]. In contrast, our work investigates the convergence of sPPA with a large constant step size for the special case of a least-squares loss. Such an approach allows for aggressive updates throughout the algorithm, potentially improving its practical efficiency.
The ReBlocK algorithm is additionally a special case of the SlimLS algorithm proposed by [chung2020sampled] for solving massive linear inverse problems. In fact, their convergence result for the expectation of the ReBlocK iterates, found in result (a) of their Theorem 3.1, is equivalent to equation (25) of our Theorem 5.1. However, by identifying the limiting expectation value as a weighted least-squares solution, we provide greater conceptual clarity regarding its meaning. Additionally, our bounds regarding the variance of the ReBlocK iterates, the convergence of tail averages, and the suboptimality of the weighted least-squares solution are all much stronger than the corresponding bounds of [chung2020sampled]. These stronger bounds serve to explain our numerical findings that ReBlocK can be robust and efficient even when $\lambda$ is very small and the problem contains many nearly singular blocks $A_{S}$, which is the most relevant case and is not explained by the previous analysis.
Our work is also related to the nearly concurrent paper [derezinski2025], which introduces Tikhonov regularization into the RBK iterations just like ReBlocK. [derezinski2025] focuses on consistent systems and in this context, the regularization gives rise to optimal convergence rates in the presence of Nesterov acceleration. On the other hand, our work focuses on solving inconsistent systems, in which case the regularization is needed to ensure the stability of the algorithm. Exploring the combination of regularization and Nesterov acceleration in the inconsistent case is a promising direction for future work.
@@ 2 Overview of the Analysis
Here $M(\cdot):\mathbb{R}^{k\times n}\rightarrow\mathbb{R}^{k\times k}$ takes in the sampled block $A_{S_{t}}$ and returns a positive semidefinite “mass” matrix. RBK is recovered by setting $M(A_{S_{t}})=(A_{S_{t}}A_{S_{t}}^{\top})^{+}$ and ReBlocK by setting $M(A_{S_{t}})=(A_{S_{t}}A_{S_{t}}^{\top}+\lambda kI)^{-1}$. See Algorithm 1 for the full procedure with and without tail averaging.
@@
where $I_{S}$ represents the $k$ rows of the $n\times n$ identity matrix whose indices are in $S$ (recall $|S|=k$). These quantities are natural because they enable the general iteration (8) to be rewritten as
@@
Since the normal equations for (12) can be written as $A^{\top}\overline{W}r^{(\rho)}=0$, we observe that the final term in (13) vanishes in expectation. Indeed, identifying the appropriate weighted solution $x^{(\rho)}$ to enable the generalized iteration (8) to be written in the form of (13), namely as a linear contraction of the error plus a zero-mean additive term, is the main technical innovation underlying our results. From here, the analysis of [epperly2024randomizedkaczmarztailaveraging] can be readily generalized to show that convergence to $x^{(\rho)}$ is obtained.
@@ 3 RBK without Preprocessing
We begin by presenting our general convergence bounds for RBK-U. We emphasize that these bounds should not be viewed as a proof that algorithm is effective, but rather as a helpful framework to understand the properties and the potential problems of the algorithm. Indeed we will later demonstrate severe failure modes of RBK-U that are inspired by the analysis (Equation 18, Figure 2) and we will ultimately show that regularization is needed in order to eliminate these failure modes and obtain satisfying bounds on the bias and the variance (Theorem 5.1).
@@
(15) $\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq\left(1-\alpha\right)^{T_{b}+1}\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}+\frac{2}{\alpha^{2}(T-T_{b})}\;V$
with $V=\mathbb{E}\,_{S\sim\rho}\|A_{S}^{+}r^{(\rho)}_{S}\|^{2}$. Finally, when the rows of $A$ are in general position the condition number $\kappa(\overline{W})$, the weighted residual $r^{(\rho)}$, the bias $x^{(\rho)}-x^{*}$, and the variance $V$ are bounded as
(16) $\kappa(\overline{W})$ $\leq k\cdot\max_{i}\left\lVert a_{i}\right\rVert^{2}\cdot\max_{S}\|A_{S}^{+}\|^{2},$
$\|r^{(\rho)}\|$ $\leq\sqrt{\kappa(\overline{W})}\cdot\left\lVert r\right\rVert,$
$\|x^{(\rho)}-x^{*}\|$ $\leq\sqrt{\kappa(\overline{W})-1}\cdot\left\lVert A^{+}\right\rVert\cdot\left\lVert r\right\rVert,$
$V$ $\leq\max_{S}\|A_{S}^{+}\|^{2}\cdot\kappa(\overline{W})\cdot\frac{k\left\lVert r\right\rVert^{2}}{m}.$
@@
To our knowledge, this is the first result for inconsistent linear systems that characterizes the exact solution to which the randomized block Kaczmarz iterates converge, albeit in a Monte Carlo sense. The $O(1/T)$ convergence rate for the tail-averaged bound is optimal for row-access methods, and a reasonable default for the burn-in time is $T_{b}=T/2$; see [epperly2024randomizedkaczmarztailaveraging] for a more thorough discussion of these points. The proof of Theorem 3.1, which takes advantage of an orthogonal decomposition of the error term $x_{t+1}-x^{(\rho)}$, can be found in Appendix B.
Unfortunately, Theorem 3.1 does not imply robust convergence for general problems. Indeed, for problems containing nearly singular blocks $A_{S}$, the bounds on the condition number, the weighted residual, the bias, and the variance can be arbitrarily large. In fact, these issues are not just defects of the analysis but real problems with RBK-U. To substantiate this, we present a single-parameter family of $3\times 2$ examples that completely breaks RBK-U:
Example 3.2 (No-go for RBK-U).
Let $\epsilon>0$ and
(17) $A=\begin{bmatrix}0&1\\ 1&\epsilon^{2}\\ 1&-\epsilon^{2}\end{bmatrix},\quad b=\begin{bmatrix}0\\ 1+\epsilon\\ 1-\epsilon\end{bmatrix}$.
Then $\lim_{\epsilon\rightarrow 0^{+}}\kappa(A)=\sqrt{2}$ and $\lim_{\epsilon\rightarrow 0^{+}}\left\lVert r\right\rVert=0$, indicating that the problem is well conditioned and approaches consistency as $\epsilon\rightarrow 0$. On the other hand, as $\epsilon\rightarrow 0$ the bottom $2\times 2$ block approaches singularity and so for $k=2$ the upper bounds of $\kappa(\overline{W})$, $\left\lVert r^{(\rho)}\right\rVert$, $\left\lVert x^{(\rho)}-x^{*}\right\rVert$, and $V$ from Theorem 3.1 all diverge in this limit. In fact, this divergence is necessary as it can be verified that for RBK-U with $k=2$, it holds
(18) $\lim_{\epsilon\rightarrow 0^{+}}\kappa(\overline{W})=\lim_{\epsilon\rightarrow 0^{+}}\|r^{(\rho)}\|=\lim_{\epsilon\rightarrow 0^{+}}\|x^{(\rho)}-x^{*}\|=\lim_{\epsilon\rightarrow 0^{+}}V=\infty.$
The fact that the bias and variance of the RBK-U iterates diverges while the underlying problem is approaching consistency represents a clear and catastrophic failure of the algorithm. We can understand this failure by noting that geometrically, this problem represents three constraints in two dimensions which together form an isosceles triangle with vertices at $(1-\epsilon,0)$, $(1+\epsilon,0)$, and $(1,1/\epsilon)$. When running RBK-U with $k=2$, the algorithm jumps between the three vertices of the triangle with equal probabilities, and as a result $x^{(\rho)}$ coincides with the centroid of the triangle. The problem is that as $\epsilon\rightarrow 0$ the top vertex walks off to infinity and drags the centroid with it, while the least-squares solution $x^{*}$ converges to $(1,0)$. This insight can be used to directly show the divergence of the bias, the residual, and the variance. The residual bound $\|r^{(\rho)}\|\leq\sqrt{\kappa(\overline{W})}\cdot\left\lVert r\right\rVert$ then implies the divergence of $\kappa(\overline{W})$ since $\|r^{(\rho)}\|$ diverges while $\left\lVert r\right\rVert$ goes to zero. The situation is depicted visually in Fig. 1.
$\bm{a}_{1}^{\top}\bm{x}=b_{1}$$\bm{a}_{2}^{\top}\bm{x}=b_{2}$$\bm{a}_{3}^{\top}\bm{x}=b_{3}$$x^{\ast}$$x^{(\rho)}$
Figure 1: Visual depiction of the $3\times 2$ linear system Eq. 17 that causes RBK-U to fail catastrophically when $k=2$. As $\epsilon\rightarrow 0$ the top vertex walks off to infinity and the recovered solution $x^{(\rho)}$ goes with it, while the true solution $x^{*}$ approaches the $x$-axis.
It is worth noting that such problems are not limited to any particular batch size such as $k=2$. More generally we can view the skinny isosceles triangle as a stretched $2$-simplex and indeed, the same arguments can be readily generalized to show that for arbitrary $k$, a similarly stretched $k$-simplex suffices to break RBK-U with a block size of $k$. The problem is also not directly related to the use of uniform sampling, and similar problems could arise for any a priori sampling distribution.
3.1 Numerical demonstration
We now demonstrate that the issues highlighted by Equation 18 can occur for more realistic problems. To do so, we set $m=10^{5}$, $n=10^{2}$ and construct two inconsistent problems whose columns are discretized representations of continuous functions, leading the matrices A to contain many nearly singular blocks of rows. We apply both minibatch SGD (mSGD) and RBK-U to the resulting problems, applying tail averaging to each algorithm to observe convergence beyond the variance horizon. The results in Fig. 2 confirm that RBK-U performs poorly for these problems, just as our theory suggests. In fact RBK-U is not even able to reliably attain a relative error of $1.0$ in either case, meaning the algorithm performs worse than simply guessing $x^{*}=0$. For these and other experiments in Sections 3, 4, and 5 we choose the burn-in time $T_{b}$ to demonstrate the different phases of convergence as clearly as possible.
We note that for such realistic examples, it is difficult to pin down exactly how large the bias $x^{(\rho)}-x^{*}$ is. The reason is that the variance $V$ is so large that running the algorithm to convergence would be prohibitively expensive. Nonetheless it is evident that RBK-U is not effective for these problems, even after incorporating tail averaging. The code for all of our experiments can be found at https://github.com/ggoldsh/block-kaczmarz-without-preprocessing, and further details on these particular experiments can found in Appendix F.
Figure 2: Failure of RBK-U for two inconsistent problems with many nearly singular blocks, with mild singular value decay (left) and rapid singular value decay (right). The vertical dotted line indicates the burn-in time, before which results are shown for individual iterates.
3.2 Implementation Details
To stably implement the RBK iteration (5), we employ a QR-based least-squares solver to calculate $A_{S_{t}}^{+}(b_{S_{t}}-A_{S_{t}}x).$ The most expensive part of this procedure is the QR decomposition, which has an asymptotic cost of $O(nk^{2})$. This cost is greater than the cost of an mSGD iteration, which is $O(nk)$.
4 Linear Least-squares with Gaussian Data
Given the problems with RBK-U for general data, it is natural to wonder if there are more restricted categories of benign problems for which the bias and variance of RBK-U can be controlled. Previous works such as [derezinski2024fine, derezinski2024solving, derezinski2025] have characterized the performance of RBK-U for consistent systems that have been preprocessed with a randomized Hadamard transform. After preprocessing the resulting system is highly unlikely to contain nearly singular blocks or other problematic coherence properties, leading to strong convergence guarantees for RBK-U. However, it is not clear why matrices that have not been preprocessed would resemble such preprocessed matrices. Thus, instead of directly extending these results to the inconsistent case, we instead consider inconsistent problems whose rows arise from a Gaussian distribution, which seem more likely to appear in practice. Furthermore, to simplify the analysis we consider the infinite version of the problem in which each uniformly selected row is drawn directly from the corresponding Gaussian distribution. This enables us to exploit an elegant connection between uniform block samples of Gaussian data and dense Gaussian sketches of arbitrary data. It should be possible to extend our results to provide high probability bounds for finite problems using more involved techniques such as those of [derezinski2024fine, derezinski2024solving, derezinski2025], but we consider this beyond the scope of the current work.
Now, consider the statistical least-squares problem
(19) $x^{*}=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\mathbb{E}_{[a^{\top}\;b]\sim\mathcal{N}(0,Q)}\left[(a^{\top}x-b)^{2}\right].$